The radioactive nuclides Pt - 199 has a half-life of 30.8 minutes. A sample is prepared that has an initial activity of 7.56 x 10^11Bq. a) How many Pt -199 nuclei are initially present in the sample?
b) How many are present after 30.8 minutes?
c) What is the activity at this time?
1
Expert's answer
2020-11-12T09:16:05-0500

a)


"N_0=\frac{A}{\lambda}=\frac{7.56 \cdot 10^{11}(30.8\cdot60)}{\ln{2}}=2.02\cdot 10^{15}"

b)


"N=0.5^{\frac{\tau}{\tau}}N=0.5(2.02\cdot 10^{15})=1.01\cdot 10^{15}"

c)


"A(t)=0.5^{\frac{\tau}{\tau}}A=0.5(7.56 \cdot 10^{11})=3.78 \cdot 10^{11}\ Bq"